# On the controller form for linear hyperbolic MIMO systems with dynamic boundary conditions

### v1

@@ Abstract
This contribution develops an algebraic approach to obtain a controller form for a class of linear hyperbolic MIMO systems, bidirectionally coupled with a linear ODE system at the unactuated boundary. After a short summary of established controller forms for SISO and MIMO ODE as well as SISO hyperbolic PDE systems, it is shown that the direct approach to state a controller form fails already for a very simple MIMO example. Next, a generalised hyperbolic controller form with different variants is proposed and a new flatnesss-based scheme to compute said form is presented. Therein, the system is treated in an algebraic setting where generalised polynomials with real exponents are used to describe the predictions and delays in the system. The proposed algorithm is then applied to the motivating example.
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
From the onset of state space based approaches, controller (canonical) forms have been a vital asset for systematic controller designs. As the controller forms for linear SISO and MIMO ODE systems, cf. [1, 2] as well for SISO hyperbolic PDE systems [3] are long established and understood, only comparatively recent works [4, 5, 6] discuss hyperbolic systems with boundaries governed by ODEs for the SISO case which this contribution extends to the MIMO case.
Following some preliminaries, where we briefly restate established controller forms and demonstrate the problems that arise if a hyperbolic controller form is to be introduced even for the most simple MIMO system, a MIMO hyperbolic controller form (HCF) is given in Section III. The main part of the contribution in Section IV then discusses how a HCF can be obtained for a specific class of systems whereupon Section V tracks the procedure for an example system.
@@ II-A Notation
The variables $z$ and $t$ denote space and time. The first and second derivative of $x(t)$ is denoted by $\dot{x}(t)$ and $\ddot{x}(t)$, while the $n$-th derivative is given by $x^{(n)}(t)$ and the partial derivative of $x(z,t)$ w.r.t. $z$. by $\partial_{z}x(z,t)$.
To treat delay differential equations (DDEs) in an algebraic setting, we identify the operators therein with their Fourier-Laplace transforms and treat the application of the transformed operators as convolutions. Thus, we write $\mathcal{L}\{\partial_{t}\}=s$ for differentials, while the shifts $(\sigma^{\tau}x)(t)=x(t+\tau)$ with $\tau\in\mathbb{R}$ are denoted as $\mathcal{L}\{\sigma^{\tau}\}(s)=\exp(s\tau)$.
The shorthand $L^{2}([a,b])$ is used for the space of square integrable functions on $[a,b]$, while $H^{n}([a,b])$ stands for the space of functions whose $n$-th derivative belongs to $L^{2}([a,b])$ and $\mathcal{E}^{\prime}$ for Schwartz-Distributions with compact support. With $\operatorname{V}_{f}([a,b])$ denoting the total variation of a function $f$ on the interval $[a,b]$, $BV([a,b])$ contains functions of bounded variation while members of its subset $BV^{\star}([a,b])$ satisfy $\lim_{\epsilon\to 0}\operatorname{V}_{f}([b,b+\epsilon))=0$.
Furthermore, $\mathbb{R}[\partial_{t}]$ is the space of polynomials in $\partial_{t}$ with coefficients in $\mathbb{R}$ and $\mathbb{R}(\partial_{t})$ consists of rational expressions in $\partial_{t}$. Regarding a polynomial matrix $X(\partial_{t})\in\mathbb{R}^{n\times n}[\partial_{t}]$, we denote the leading column coefficient matrix (LCCM) w.r.t. $\partial_{t}$ by $\operatorname{lccm}_{\partial_{t}}{\mathbf{X}(\partial_{t})}$ and say that $\mathbf{X}(\partial_{t})$ is column reduced w.r.t. $\partial_{t}$ if $\operatorname{lccm}_{\partial_{t}}{\mathbf{X}(\partial_{t})}$ has full generic rank. Finally, we call $\mathbf{X}(\partial_{t})$ unimodular if its determinant is independent of $\partial_{t}$.
@@ II-B1 ODE SISO systems
with state $\bm{\eta}^{\mathrm{O}}(t)=(\eta^{\mathrm{O}}_{1}(t),\dotsc,\eta^{\mathrm{O}}_{n}(t))^{\intercal}\in\mathbb{R}^{n}$, coefficients $\bm{a}\in\mathbb{R}^{n}$ and input $u(t)\in\mathbb{R}$ that exists for every controllable system. Note that the “output” $\eta^{\mathrm{O}}_{1}(t)$ of the integrator chain is a flat output of the system.
@@ II-B2 ODE MIMO systems
$\dot{\eta}^{\mathrm{O}}_{i,\nu_{i}}(t)$ $=-\sum_{k=1}^{m}\bm{a}_{k}^{\intercal}\bm{\eta}^{\mathrm{O}}_{k}(t)+\bm{b}_{i}^{\intercal}\bm{u}(t)$ (2c)
for $i\in\{1,\dotsc,{m}\}$, with the Kronecker index $\nu_{i}$ and the state vector $\bm{\eta}^{\mathrm{O}}_{i}(t)=(\eta^{\mathrm{O}}_{i,1}(t),\dotsc,\eta^{\mathrm{O}}_{i,\nu_{i}}(t))^{\intercal}\in\mathbb{R}^{\nu_{i}}$ of the $i$-th subsystem. Similarly to the SISO case, the MIMO controller form state $\bm{\eta}^{\mathrm{O}}(t)\in\mathbb{R}^{\nu_{\Sigma}}$ with $\nu_{\Sigma}\coloneq\sum_{i=1}^{m}\nu_{i}$ and input $\bm{u}(t)$ can be easily parametrised by the tuple $(\eta^{\mathrm{O}}_{1,1}(t),\dotsc,\eta^{\mathrm{O}}_{m,1}(t))$.
II-B3 hyperbolic SISO systems
@@
$\dot{\eta}^{\mathrm{O}}_{n}(t)$ $=\eta^{\mathrm{P}}(\check{\tau},t)$ (3c)
[…]
$\eta^{\mathrm{P}}(\hat{\tau},t)$ $=-\operatorname{\mathcal{S}}_{a^{\star}}\eta^{\mathrm{P}}(\cdot,t)-\bm{a}^{\intercal}\eta^{\mathrm{O}}(t)+u(t).$ (3e)
[…]
@@ II-C Benchmark System
$\partial_{t}^{2}x_{1}(z,t)-\partial_{z}^{2}x_{1}(z,t)$ $=0\quad z\in(-\alpha,0)$ (4a)
$\partial_{t}^{2}x_{2}(z,t)-\partial_{z}^{2}x_{2}(z,t)$ $=0\quad z\in(0,\beta),$ (4b)
coupled with a point mass at $z=0$ and thus subject to the dynamic boundary conditions
@@
$\partial_{z}x_{1}(-\alpha,t)$ $=-u_{1}(t)$ (5c)
$\partial_{z}x_{2}(\beta,t)$ $=u_{2}(t),$ (5d)
with state [^2] In detail, $\bm{x}(\cdot,t)\!=\!(\xi(t),\dot{\xi}(t),x_{1}(\cdot,t),\partial_{t}x_{1}(\cdot,t),x_{2}(\cdot,t),\partial_{t}x_{2}(\cdot,t))^{\intercal}$ and $\mathcal{X}=\{\mathbb{R}^{2}\times L^{2}([-\alpha,0])\times H^{1}([-\alpha,0])\times L^{2}([0,\beta])\times H^{1}([0,\beta])\;|\;\xi(t)=x_{1}(0,t)=x_{2}(0,t)\}$. $\bm{x}(\cdot,t)\in\mathcal{X}$ as initially discussed in [7]. By using the deflection and gradient at $z=0$ as Cauchy data, D’Alemberts’s formula yields the solutions
$x_{i}(z,t)\!=\!\frac{1}{2}\!\left(x_{i}(0,t-z)\!+\!x_{i}(0,t+z)\right)\!+\!\frac{1}{2}\!\int\limits_{t-z}^{t+z}\partial_{z}x_{i}(0,\tau)\,\mathrm{d}\tau$ (6)
for $i=1,2$ (c.f. [7, 4]), which reads
$x_{i}(z,t)=C_{z}x_{i}(0,t)+\frac{1}{s}S_{z}\partial_{z}x_{i}(0,t)$ (7)
in the algebraic setting, where $C_{z}\coloneqq\cosh(\partial_{t}z)=\frac{1}{2}(\sigma^{z}+\sigma^{-z})$ and $S_{z}\coloneqq\sinh(\partial_{t}z)=\frac{1}{2}(\sigma^{z}-\sigma^{-z})$.
II-D Parametrisation
@@
with $y_{1}\in H^{2}(\mathbb{R})$ and $y_{2}\in L^{2}(\mathbb{R})$ we directly obtain $\xi(t)=x_{1}(0,t)=x_{2}(0,t)={y}_{1}(t)$ as well as
$\partial_{z}x_{1}(0,t)$ $=\frac{1}{2}\left({y}_{2}(t)-m\partial_{t}^{2}{y}_{1}(t)\right)$ (9a)
$\partial_{z}x_{2}(0,t)$ $=\frac{1}{2}\left({y}_{2}(t)+m\partial_{t}^{2}{y}_{1}(t)\right)$ (9b)
from (5b). With the Cauchy data known, (7) then yields the profiles, from which the input parameterisation
$\begin{pmatrix}u_{1}(t)\\ u_{2}(t)\\ \end{pmatrix}=\begin{pmatrix}\frac{m}{2}\partial_{t}^{2}C_{\alpha}+\partial_{t}S_{\alpha}&-\frac{1}{2}C_{\alpha}\\ \frac{m}{2}\partial_{t}^{2}C_{\beta}+\partial_{t}S_{\beta}&\frac{1}{2}C_{\beta}\end{pmatrix}\begin{pmatrix}y_{1}(t)\\ y_{2}(t)\\ \end{pmatrix}$ (10)
@@ II-E Stating a controller form
As known from, e.g [4], the problem to be solved for stating a controller form is to find a particular realisation of the input-output relation (10). A possible way to do so is to solve (10) for the highest derivatives of the largest predictions of $y_{1}(t)$ as well as $y_{2}(t)$ and to check for which initial conditions the resulting DDE system is well posed.
However, taking the coefficients of the highest derivatives of the largest predictions in each column on the right hand-side of (10), the resulting leading column coefficient matrix
$\operatorname{lccm}_{\rho,s}{\mathbf{H}}=\frac{1}{4}\begin{pmatrix}0&0\\ m&1\\ \end{pmatrix}$ (11)
does not have full rank. This implies that (10) cannot directly be written as an explicit DDE system thus negating our approach. Therefore, more steps have to be taken for MIMO systems, which will be discussed in Section IV after we give a definition of the desired controller form for this case.

### v3

@@ Abstract
This contribution develops an algebraic approach to obtain a controller form for a class of linear hyperbolic MIMO systems, bidirectionally coupled with a linear ODE system at the unactuated boundary. After a short summary of established controller forms for SISO and MIMO ODE as well as SISO hyperbolic PDE systems, it is shown that the approach to state a controller form for SISO systems cannot easily be transferred to the MIMO case as it already fails for a very simple example. Next, a generalised hyperbolic controller form with different variants is proposed and a new flatness-based scheme to compute said form is presented. Therein, the system is treated in an algebraic setting where quasipolynomials are used to express the predictions and delays in the system. The proposed algorithm is then applied to the motivating example.
©2026 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
Canonical forms serve as a cornerstone for systematic feedback design in finite dimensions. Their counterparts for distributed parameter systems (DPSs) have been studied since the late 1970s. For linear hyperbolic partial differential equations (PDEs), [1] connected controllability and spectral assignment to controller canonical forms (CCFs) and neutral delay equations. More recent works linked the aforementioned delay interpretations to flatness based parametrisations [2, 3] (as well as [4] for early ideas in this direction) and established the connection between such parametrisations and the CCF for more general systems. In particular, [5] introduced a hyperbolic controller canonical form (HCCF) for linear hyperbolic SISO systems comprising general interconnections of hyperbolic PDEs and ordinary differential equations (ODEs) (see also [6]) by generalising the familiar chain-of-integrators structure by prepending a transport equation to represent the delays of the system imposed by the hyperbolic PDEs. Therein, the canonical coordinates are deduced from the feasible initial conditions (ICs) for a functional differential equation (FDE) arising from a flatness-based parametrisation of the system’s input while the required state transforms are computed using the flatness-based parametrisation of the system’s original state variables. Aside from the above-mentioned state-space approaches, in [7, 8] flatness-based parametrisations also served as means for purely algebraic analysis and control design. These methods employ and partially generalise results from [9] and [10].
As many relevant technological applications have multiple in- and outputs, several control approaches are available, e.g.  [11] using flatness-based methods or [12] employing backstepping designs. However, regarding controller forms a direct extension of the existing work to the MIMO case is difficult, as such systems can exhibit different delays depending on the considered input which together with the bidirectionally coupled ODE result in an FDE system for which feasible ICs are not obvious, thus, negating the approach from the SISO case.
Therefore, the goal of this contribution is to introduce a hyperbolic controller form (HCF) for a class of linear hyperbolic MIMO systems with dynamic boundary conditions and to develop a delay reduction algorithm such that the resulting FDE allows for a direct computation of the controller form. To do so, we denote the input-output relation using quasipolynomials, whose use in the control of hyperbolic systems was already discussed in [13, 14], and restate the requirements to introduce a HCF as column-reduced properties on the resulting quasipolynomial matrices. Next, in contrast to the established methods and due to the incommensurate delays in the considered systems as well as the interplay with the boundary system, we propose a modified quasipolynomial division algorithm to compute the required transformation to establish said properties.
Following some preliminaries, where we briefly restate established controller forms and demonstrate the problems that arise if a hyperbolic controller form is to be introduced even for the most simple MIMO system, a general MIMO HCF and several subforms are given in Section III. The main part of the contribution in Section IV then discusses how a HCF can be obtained for a specific class of systems Finally, Section V traces the procedure for an example system.
@@ II-A Notation
The variables $z$ and $t$ denote space and time. The first and second derivative of $x(t)$ is denoted by $\dot{x}(t)$ and $\ddot{x}(t)$, while the $n$-th derivative is given by $x^{(n)}(t)$ and the partial derivative of $x(z,t)$ w.r.t. $z$ by $\partial_{z}x(z,t)$.
To treat FDEs in an algebraic setting, we identify the operators therein with their Fourier-Laplace transforms and treat the application of the transformed operators as convolutions. Thus, we write $\mathcal{L}\{{\mathrm{d}_{t}}\}=s$ for differentials, while the shifts $(\sigma^{\tau}x)(t)=x(t+\tau)$ with $\tau\in\mathbb{R}$ are denoted as $\mathcal{L}\{\sigma^{\tau}\}(s)=\exp(s\tau)$.
The shorthand $L^{2}([a,b])$ is used for the space of square integrable functions on $[a,b]$, while $H^{n}([a,b])$ stands for the space of functions whose $n$-th derivative belongs to $L^{2}([a,b])$ and $\mathcal{E}^{\prime}$ for Schwartz-Distributions with compact support. With $V_{f}([a,b])$ denoting the total variation of a function $f$ on the interval $[a,b]$, $\mathrm{BV}([a,b])$ contains functions of bounded variation while members of its subset $\mathrm{BV}^{\star}([a,b])$ satisfy $\lim_{\epsilon\to 0}V_{f}([b,b-\epsilon))=0$.
Furthermore, $\mathbb{R}[{\mathrm{d}_{t}}]$ is the space of polynomials in $\mathrm{d}_{t}$} with coefficients in $\mathbb{R}$ while $\mathbb{R}({\mathrm{d}_{t}})$ consists of rational expressions in $\mathrm{d}_{t}$}. Regarding a polynomial matrix $\mathbf{X}({\mathrm{d}_{t}})\in\mathbb{R}^{n\times n}[{\mathrm{d}_{t}}]$, we denote the leading column coefficient matrix (LCCM) w.r.t. $\mathrm{d}_{t}$} by $\operatorname{lccm}_{{\mathrm{d}_{t}}}{\mathbf{X}({\mathrm{d}_{t}})}$ and say that $\mathbf{X}({\mathrm{d}_{t}})$ is column reduced w.r.t. $\mathrm{d}_{t}$} if $\operatorname{lccm}_{{\mathrm{d}_{t}}}{\mathbf{X}({\mathrm{d}_{t}})}$ has full generic rank. Finally, we call $\mathbf{X}({\mathrm{d}_{t}})$ unimodular over $\mathbb{R}[{\mathrm{d}_{t}}]$ if its determinant is a nonzero constant (and therefore independent of $\mathrm{d}_{t}$}).
@@ II-B1 ODE SISO systems
with state $\bm{\eta}^{\mathrm{O}}(t)=(\eta^{\mathrm{O}}_{1}(t),\dotsc,\eta^{\mathrm{O}}_{n}(t))^{\intercal}\in\mathbb{R}^{n}$, coefficients $\bm{a}\in\mathbb{R}^{n}$ and input $u(t)\in\mathbb{R}$, which exists for every controllable system. Note that the “output” $\eta^{\mathrm{O}}_{1}(t)$ of the integrator chain is a flat output of the system.
@@ II-B2 ODE MIMO systems
$\dot{\eta}^{\mathrm{O}}_{i,\nu_{i}}(t)$ $=-\sum_{j=1}^{m}\bm{a}_{ij}^{\intercal}\bm{\eta}^{\mathrm{O}}_{j}(t)+\bm{b}_{i}^{\intercal}\bm{u}(t)$ (2c)
for $i\in\{1,\dotsc,{m}\}$, with the Kronecker indices $\nu_{i}$ and the state vector $\bm{\eta}^{\mathrm{O}}_{i}(t)=(\eta^{\mathrm{O}}_{i,1}(t),\dotsc,\eta^{\mathrm{O}}_{i,\nu_{i}}(t))^{\intercal}\in\mathbb{R}^{\nu_{i}}$ of the $i$-th subsystem. Similarly to the SISO case, the MIMO controller form state $\bm{\eta}^{\mathrm{O}}(t)\in\mathbb{R}^{\nu_{\Sigma}}$ with $\nu_{\Sigma}\coloneq\sum_{i=1}^{m}\nu_{i}$ and input $\bm{u}(t)$ can be easily parametrised by a flat output $(\eta^{\mathrm{O}}_{1,1}(t),\dotsc,\eta^{\mathrm{O}}_{m,1}(t))$.
Note that, in principle, there exist several ways to distribute the $\nu_{\Sigma}$ integrators over the m{m} chains but only one that yields the form (2) without input derivatives.
II-B3 Hyperbolic SISO systems with boundary dynamics
@@
$\dot{\eta}^{\mathrm{O}}_{n}(t)$ $=\eta^{\mathrm{P}}(0,t)$ (3c)
$\partial_{t}\eta^{\mathrm{P}}(\tau,t)$ $=\partial_{\tau}\eta^{\mathrm{P}}(\tau,t),\quad\tau\in[0,\hat{}\tau)$ (3d)
$\eta^{\mathrm{P}}(\hat{}\tau,t)$ $=-\left(a^{\star}\eta^{\mathrm{P}}\right)(t)-\bm{a}^{\intercal}\eta^{\mathrm{O}}(t)+u(t)$ (3e)
with state $\bm{\eta}(\cdot,t)=(\bm{\eta}^{\mathrm{O}}(t),\eta^{\mathrm{P}}(\cdot,t))\in\mathbb{R}^{n}\times L^{2}([0,\hat{}\tau])$ and the Stieltjes integral $\left(a^{\star}\eta^{\mathrm{P}}\right)(t)=\int_{0}^{\hat{}\tau}\eta^{\mathrm{P}}(\tau,t)\,\mathrm{d}\bar{a}(\tau)$ where $\bar{a}\in BV^{\star}([0,\hat{}\tau])$ allows for distributed feedback of $\eta^{\mathrm{P}}(z,t)$ as well as point evaluations for $z\in[0,\hat{}\tau)$. Similar to the SISO case, $\eta^{\mathrm{O}}_{1}(t)$ is a flat output of the system.
@@ II-C Benchmark system
$\partial_{t}^{2}x_{1}(z,t)-\partial_{z}^{2}x_{1}(z,t)$ $=0\quad z\in(-{l_{1}},0)$ (4a)
$\partial_{t}^{2}x_{2}(z,t)-\partial_{z}^{2}x_{2}(z,t)$ $=0\quad z\in(0,{l_{2}}),$ (4b)
coupled with a point mass M{M} at $z=0$ and thus subject to the dynamic boundary conditions
@@
$\partial_{z}x_{1}(-{l_{1}},t)$ $=-u_{1}(t)$ (5c)
$\partial_{z}x_{2}({l_{2}},t)$ $=u_{2}(t),$ (5d)
with state [^2] In detail, $\bm{x}(\cdot,t)\!=\!(\xi(t),\dot{\xi}(t),x_{1}(\cdot,t),\partial_{t}x_{1}(\cdot,t),x_{2}(\cdot,t),\partial_{t}x_{2}(\cdot,t))^{\intercal}$ and $\mathcal{X}=\{\mathbb{R}^{2}\times H^{1}([-{l_{1}},0])\times L^{2}([-{l_{1}},0])\times H^{1}([0,{l_{2}}])\times L^{2}([0,{l_{2}}])\;|\;\xi(t)=x_{1}(0,t)=x_{2}(0,t)\}$. $\bm{x}(\cdot,t)\in\mathcal{X}$ as initially discussed in [18].
II-D Input parametrisation by a flat output
@@
with $y_{1}\in H^{2}(\mathbb{R})$ as well as $y_{2}\in L^{2}(\mathbb{R})$ we directly obtain $\xi(t)=x_{1}(0,t)=x_{2}(0,t)={y}_{1}(t)$ and via (5b)
$\partial_{z}x_{1}(0,t)$ $=\frac{1}{2}\left({y}_{2}(t)-{M}{\mathrm{d}_{t}}^{2}{y}_{1}(t)\right)$ (7a)
$\partial_{z}x_{2}(0,t)$ $=\frac{1}{2}\left({y}_{2}(t)+{M}{\mathrm{d}_{t}}^{2}{y}_{1}(t)\right).$ (7b)
By using the deflection $x_{i}(0,t)$ and its gradient $\partial_{z}x_{i}(0,t)$ as Cauchy data, the Method of Characteristics yields the profiles
$x_{i}(z,t)\!=\!\frac{1}{2}\!\left(\!\xi(t-z)\!+\!\xi(t+z)\!+\!\!\int_{t-z}^{t+z}\!\partial_{z}x_{i}(0,\tau)\,\mathrm{d}\tau\right)$ (8)
for $i=1,2$ (cf. [18, 5]), which can be written as
$x_{i}(z,t)=C_{z}x_{i}(0,t)+S_{z}\partial_{z}x_{i}(0,t)$ (9)
in the algebraic setting, where $C_{z}=\frac{1}{2}(\sigma^{z}+\sigma^{-z})$ and $S_{z}=\frac{1}{2}\frac{\sigma^{z}-\sigma^{-z}}{{\mathrm{d}_{t}}}$ employ $\sigma$ as well as $\mathrm{d}_{t}$} for shifts and derivatives.
Herein, note that although $S_{z}$ contains a division by $\mathrm{d}_{t}$} which results in an integration in the parametrisation of $x_{i}(z,t)$ by $\bm{y}(t)$, this does not contradict the idea of a unique parametrisation as $S_{z}\in\mathcal{E}^{\prime}$ has compact support and the corresponding integral in (8) therefore is definite.
With the parametrisation of the Cauchy data from (7), (9) then yields the profiles, thus, the input parametrisation
$\begin{pmatrix}u_{1}(t)\\ u_{2}(t)\\ \end{pmatrix}=\begin{pmatrix}\frac{m}{2}{\mathrm{d}_{t}}^{2}C_{l_{1}}+{\mathrm{d}_{t}}S_{l_{1}}&-\frac{1}{2}C_{l_{1}}\\ \frac{m}{2}{\mathrm{d}_{t}}^{2}C_{l_{2}}+{\mathrm{d}_{t}}S_{l_{2}}&\frac{1}{2}C_{l_{2}}\end{pmatrix}\begin{pmatrix}y_{1}(t)\\ y_{2}(t)\\ \end{pmatrix}$ (10)
@@ II-E Stating a controller form
As known from e.g. [5], the problem to be solved for stating a controller form is to find a particular realisation of the input-output relation (10). A possible way to do so is to solve (10) for the highest derivatives of the largest predictions of $y_{1}(t)$ as well as $y_{2}(t)$ and to check for which initial conditions the resulting FDE system is well-posed.
However, the leading column coefficient matrix
$\hat{}\mathbf{G}=\frac{1}{4}\begin{pmatrix}0&0\\ {M}&1\\ \end{pmatrix}$, (11)
obtained by taking the coefficients of the highest derivatives of the largest predictions in each column on the right hand-side of (10), is singular. This implies that (10) cannot directly be written as an explicit FDE system thus negating our approach. Therefore, more steps must be taken in the MIMO case, which will be discussed in Section IV after we give a definition of a controller form for this system class.